Three persons are walking from a place A to another place B. Their speeds are in the ratio of 4 : 3 : 5. The time ratio to reach B by these persons will be :


[A] 4 : 3 : 5


[B] 5 : 3 : 4


[C] 15 : 9 : 20


[D] 15 : 20 : 12 

Grade:9

2 Answers

jhansi adapaka
10 years ago

Ans: D

we need ratio''s time

time = distance/velocity

let the distance be x

and they have given ratio''s of velocities as 4:3:5

Now the ratio''s of time to reach B from A is x/4:x/3:x/5

1/4:1/3:1/5

15:20:12

Khandavalli Satya Srikanth
10 years ago

The answer for this question is D.

We know that time=distance/speed.Three persons walk same distance.Therefore time=1/speed.

(1/4):(1/3):(1/5)=(60/4):(60/3):(60/5)=15:20:12